2e^x is a particular integral for y'-y=2e^x. Is it true or false. Give reasons for your answer


1
Expert's answer
2021-02-24T07:35:55-0500

Here we have, the differential equation "y'-y=2e^x" where "y'=\frac{dy}{dx}"


Now, the co-efficient of "y" is "-1" .


So, let "\mu(x)=e^{\int-1dx}=e^{-x}"


Now, multiplying both sides of the differential equation by "\mu(x)" we get:-



"\mu(x)y'-\mu(x)y=2e^x\mu(x)\\
\Rightarrow e^{-x}\cdot\frac{dy}{dx}-ye^{-x}=2\\
\Rightarrow e^{-x}\cdot \frac{d}{dx}(y)+y\cdot\frac{d}{dx}(e^{-x})=2\\
\Rightarrow \frac{d}{dx}(ye^{-x})=2\\
\Rightarrow d(ye^{-x})=2dx\\"

Integrating both sides we have:-



"\int d(ye^{-x})=2\int dx+C \>\>\>\> (C(constant)\in\R)\\
\Rightarrow ye^{-x}=2x+C"

Now, to check if "y=2e^x" is a solution we substitute "y" in place of the general solution:-


"2e^x\cdot e^{-x}=2x+C\\
\Rightarrow C=2-2x"


As "C" is not a constant (i.e. independent of x) so "y=2e^x" is not a solution for the given differential equation.